# Planning and Rendering: Towards End-to-End Product Poster Generation

### v1

@@ 6 Conclusion
We propose a novel product poster generation framework named P&R, which can generate product posters with well-designed layouts and visually appealing backgrounds. In P&R, a PlanNet is firstly used to plan the overall layout, which merges the product image and text contents. This facilitates the creation of a well-organized layout. Then a RenderNet employs a spatial fusion module to incorporate the layouts generated by PlanNet, and encodes the appearance of the product to ensure a harmonious blend between the generated background and the product image. To promote research in this field, we construct the PPG30k dataset for end-to-end product poster generation task. Extensive experiments on PPG30k demonstrate that P&R surpasses state-of-the-art methods in terms of layout prediction and poster quality. In the future, we will focus on integrating the visual text generation model into a unified framework.

### v2

@@ 6 Conclusion
In this paper, we propose a novel Product Poster Generation (PPG) task and a corresponding product poster generation framework with diffusion models named P&R, which can generate product posters with well-designed layouts and visually appealing backgrounds. In P&R, a PlanNet is firstly used to plan the overall layout, which merges the product image and text contents. This facilitates the creation of a well-organized layout. Then a RenderNet employs a spatial fusion module to incorporate the layouts generated by PlanNet, and encodes the appearance of the product to ensure a harmonious blend between the generated background and the product image. To promote research in this field, we construct the PPG30k dataset for PPG task. Extensive experiments on PPG30k demonstrate that P&R surpasses state-of-the-art methods in terms of layout prediction and poster quality. In the future, we will focus on integrating the visual text generation model into a unified framework.